in a bomb calorimeter, 50g of a fuel was placed into the sample dish and was allowed to undergo combustion. after the combustion, the water in the calorimeter experienced an increase in temperature from 28°c to 32°c. calculate the heat of combustion of the fuel. assume the volume of the calorimeter is 4 liters and density changes in the water to be negligible. use the heat capacity of water to be 4.18 j/gºc 2.


1
Expert's answer
2021-11-08T08:27:34-0500

Q = C x m x deltaT

Cwater = 4.18 J/g x ºc

Density = m / V

The density f water is 1 g/ml. Therefore, 4 liters of water equals 4 kg.

Q = 4.18 J/g x ºc x 4000 g x (32 - 28) = 66 880 J

Heat of combustion of the fuel = Q / m = 66 880 / 50 = 1 338 J/g.